# Automated Knowledge Component Generation and Knowledge Tracing for Coding Problems

### v2

@@ 4.2 Qualitative Evaluation
Case Study Table 4 shows the estimated KC mastery levels and predicted code submission for a student on a problem in the test set. The low student mastery level on KCs “For loop iteration” and “Array indexing and assignment” results in a run time error in the predicted code by indexing the array outside of its bounds. In contrast, the higher predicted mastery level of other KCs results in the correct implementation of the if and else if statements, proper use of boolean expressions, and accurate application of the logical AND operator. This example shows that informative KC descriptions generated by the LLM can help KCGen-KT make more accurate student code predictions. In practice, the predicted mastery level may offer instructors interpretable insights into the student’s understanding of specific programming concepts.
@@
Learning Curve Analysis A common method to assess the quality of KCs is examining how well they match cognitive theory; the expected pattern on the KCs should follow the power law of practice, which states that the number of errors should decrease as the amount of practice on certain KCs increases Newell and Rosenbloom (2013); Snoddy (1926). Hence, we compare the error rate across different attempts at KCs and the estimated student KC mastery levels from KCGen-KT. For this experiment, we prompt GPT-4o to label whether each student submission contains an error on each KC. For all incorrect student submissions, we provide the problem statement, the associated KCs, and the student code, and prompt GPT-4o to (1) reason about the errors, (2) generate a corrected version of the code, and (3) assign a binary correctness label to each KC, indicating whether the student made an error on this KC in their submission (See Appendix E for the exact prompt). We acknowledge that although we find this process to be empirically accurate, a formal validation is necessary to verify the accuracy of the KC-level correctness labels.
@@
Figure 3: Representative learning curves for three generated KCs (Equality Comparison, String Length Determination, and For Loop Iteration), showing a generally decreasing error rate over attempts. Our KCs result in better model fit (0.21 vs. 0.18 in $R^{2}$) than human-written KCs under cognitive models Pavlik et al. (2009).
@@
Figure 3 shows three representative learning curves among all LLM-generated KCs. In all cases, both the ground truth and predicted error rate curves exhibit a general decreasing trend as the number of attempts increases, consistent with the power law of practice. The first predicted learning curve closely aligns with the ground truth, demonstrating KCGen-KT’s ability to accurately capture student learning progressions. The second predicted learning curve matches the ground truth error rates in trend, but has higher overall values. The third curve further exacerbates this discrepancy, for a KC that appears more frequently in the dataset. The ground truth error rates decrease overall but have significant fluctuation, making it difficult to fit the predicted learning curve. The likely reason is that students attempt problems in different orders in the CodeWorkout dataset, with some students skipping certain questions; this variation means that the same attempt may correspond to questions with different difficulties across students, making the average error rate noisy. For a more quantitative evaluation, we follow prior work Pavlik et al. (2009) and fit PFA models on each KC. Results show that the weighted $R^{2}$ metric using the LLM-generated KCs is $0.21$, and using the human-written KCs is $0.18$. Therefore, the LLM-generated KCs fit the power law of practice slightly better compared to human-written KCs.
@@
KC Ontologies We show a portion of the KC ontology subtree obtained from KCGen-KT, for string-related concepts, in Figure 4. The root node shows the KC at the highest abstraction level; going down the tree from there, we see how the KCs identified increase in granularity at each level. To build the ontology tree, we start from the top-level KC and identify all KC labels from the next level that are semantically included in it, and do it iteratively for all KCs. This mapping defines the parent-child relationships between KCs across different abstraction levels. This example demonstrates the controllable abstraction property of our KC generation pipeline, where adjusting the number of clusters directly controls the granularity of the generated KCs.
@@
Figure 4: A section of the generated KC ontology (related to Strings, at different abstraction levels.
5 Human Evaluation
We perform a human evaluation to assess the quality of the generated KCs and the accuracy of problem-KC mappings on a sample of 30 questions. We recruit 5 annotators with experience in teaching college-level programming. To evaluate KC quality, we show annotators both LLM-generated KCs and baseline human-written KCs for each problem, and ask them to indicate their preference between the two sets. On average, the LLM-generated KC sets are preferred in $54.5\%$ of the cases, suggesting that the generated KC is generally more informative to human instructors than human-written KCs.
To evaluate problem-KC mappings, we ask annotators to label every KC mapped to each problem and perform a two-stage annotation. First, they determine whether the KC is relevant to the problem. For KCs labeled as relevant, they then rate how essential it is to the problem, in three categories: essential, marginal, or non-essential. Based on these annotations, we compute the average proportion of relevant KCs per problem, as well as the average proportions of essential and non-essential KCs per problem across all annotators. Results show that the average percentage of relevant KCs per problem is $92.0\%$ for the LLM-generated set and $91.6\%$ for the baseline. The average percentage of essential KCs per problem is $50.1\%$ for the generated set and $49.5\%$ for the baseline, and the average percentage of non-essential KCs is $31.9\%$ and $33.7\%$, respectively. These findings suggest that the LLM-based KC tagging is reasonably accurate and comparable to human-labeled baselines, although there remains significant room for improvement. See Appendix D for the detailed annotation rubrics for these tasks and inter-rater agreement results.
Qualitative annotator feedback further reveals several important directions for future work. The LLM-generated KCs are generally easier to interpret and process due to their use of natural language. However, annotators note that some of these KCs could be consolidated into more concise representations, since they occasionally express overlapping concepts. Additionally, because KCs are generated independently for each problem, the LLM may overlook commonly relevant KCs present across multiple problems. Almost all feedback suggests that a human-AI collaboration approach for KC identification is a necessity: we can use KCGen-KT to provide quantitative feedback on downstream KT performance and learning curve fit, while humans merge, split, or edit LLM-generated KC descriptions and tags.
6 Conclusions and Future Work
@@
In this paper, we presented a fully automated, LLM-based pipeline for KC generation and tagging for open-ended programming problems. We also developed an LLM-based KT framework, KCGen-KT, which leverages the textual content of KC descriptions. Through extensive experiments on a real-world student coding dataset, we showed that KCGen-KT outperforms human-written KC labels on KC generation and existing state-of-the-art KT methods on predicting future student performance. We also show that LLM-generated KCs lead to better learning curves than human-written ones. A human evaluation shows that the generated KCs and problem-KC mappings are reasonably accurate to programming instructors.
@@
There are many avenues for future work. First, according to annotator suggestions, a human-in-the-loop KC generation method may address many limitations of KCGen-KT. Second, we can explore elevating the learning curve fit into the objective function and explicitly train the model to follow the power law of practice. Third, we can explore whether our methods can be applicable in other student modeling tasks and domains, including dialogues Scarlatos et al. (2025), math Ozyurt et al. (2024), and science Moore et al. (2024).
Limitations
We identify several technical and practical limitations of our work. First, the main limitation of our automatic KC generation pipeline is its reliance on in-context examples, which necessitates at least one human-written example to generate KCs at the lowest abstraction level. Without such examples, zero-shot prompting with LLMs tends to produce overly general and high-level KCs, since the LLMs are not explicitly trained for the KC generation task. Second, since each problem is associated with multiple KCs, obtaining reliable ground truth KC labels is inherently challenging. The current process of assigning ground truth correctness labels is time-consuming and requires further validation to ensure label quality. Third, the inter-rater agreement in our human evaluation is not high enough, which suggests that there is significant subjectivity in the KC evaluation task among instructors. Future work is needed to examine where this disagreement comes from and revise the evaluation process, possibly by showing KC information in alternative ways or redefining the rubrics. Fourth, we evaluate our method on a single dataset in a single domain, computer science education; applying the KC generation pipeline and KCGen-KT model across multiple datasets and domains such as math would be valuable for assessing generalizability and robustness. Finally, even though we conducted a human evaluation, the real benefit of good KC labeling is to enable students to improve learning outcomes from personalization methods informed by these KCs. Therefore, classroom studies comparing the LLM-generated KCs with human-written KCs in facilitating student progress and maximizing learning outcomes are ultimately needed.
Ethical Considerations
Our goal in this work is to develop a system that can automatically generate knowledge components and integrate them into student modeling frameworks to track individual learning progress. The primary motivation is to reduce the manual effort required from educators in topic selection and KC design, thereby enabling more time and resources to be devoted to personalized student support. However, there is a concern that such systems could replace human educator jobs, which is a shared concern across most domains with AI applications. Another critical risk lies in the quality of the automatically generated knowledge components. If the generated KCs are inaccurate, overly abstract, or misaligned with instructional goals, they may negatively affect student learning outcomes by reinforcing misconceptions or misrepresenting the required skills. Because of these reasons, we recommend that the generated knowledge component be reviewed by experts before being deployed to real students.
Appendix A Related Work
A.1 Knowledge Component Generation
Traditional methods for KC creation and tagging rely on human domain experts to identify the knowledge requirements for solving a problem Bier et al. (2014), a highly time-consuming process. Recent work has proposed automated approaches for KC discovery and tagging, employing data-driven approaches including the Q-matrix method Barnes (2005). In programming, Hosseini and Brusilovsky (2013) uses a rule-based parser to obtain ASTs with KCs identified at their lowest ontological level,  Rivers et al. (2016) define KCs as nodes in an AST followed by a learning curve analysis to identify KCs students struggle with the most in Python programming, Hoq et al. (2024) uses an AST-based neural network to identify student misconceptions, Shi et al. (2024) presents a deep learning approach for KC attribution, and Shi et al. (2023, 2024) learn latent KCs, lacking textual descriptions, by training deep learning models on KT data enforced with priors from pedagogical theory. Recent advances in LLMs have inspired automated approaches for descriptive KC generation for dialogues Scarlatos et al. (2025), and problems in math Ozyurt et al. (2024), and science Moore et al. (2024). However, we’re among the first approaches to present a fully automated, LLM-based pipeline for KC generation and tagging for open-ended programming problems.
A.2 Knowledge Tracing
There exists a wide body of work on KT Corbett and Anderson (1994) in the student modeling literature. The classic KT task aims to estimate a student’s mastery of KCs from their responses to past problems and use these estimates to predict their future performance. Classic Bayesian knowledge tracing methods Pardos and Heffernan (2010); Yudelson et al. (2013) use latent binary-valued variables to represent student KC mastery. With the widespread adoption of neural networks, multiple deep learning-based KT methods were proposed with limited interpretability since student knowledge is modeled as hidden states in these networks. Most of these methods use long short-term memory networks Hochreiter and Schmidhuber (1997) or variants Piech et al. (2015); Shin et al. (2021), with other variants coupling them with memory augmentation Zhang et al. (2017), graph neural networks Yang et al. (2020), or attention networks Ghosh et al. (2020); Pandey and Karypis (2019). KT methods have been applied to many different educational domains, including programming Hoq et al. (2023); Shi et al. (2022); Zhu et al. (2022). Recent work has attempted to leverage LLMs to develop generative KT methods predicting exact student responses to programming problems Duan et al. (2025); Fernandez and Lan (2024); Liu et al. (2022).  Sun et al. (2025) uses LLM to automatically construct Q-matrices capturing fine-grained KC relationship in knowledge tracing, while their work lacks comparison with code-specific KT baselines, and does not explore LLM-based KC generation. To the best of our knowledge, we are the first to present an LLM-based KT method for programming problems that leverages the textual content of KC descriptions, modeling interpretable student mastery levels on each KC, for improved KT performance.
Appendix B Experimental Setup
As detailed in 2.2, we use a LLM as our backbone and we use the instruction-tuned version of Llama 3 with 8B parameters as our model. We use the Parameter Efficient Fine-Tuning (PEFT) library from HuggingFace Wolf et al. (2020) load Llama 3 and fine-tune it via Low-Rank Adaptation (LoRA) Hu et al. (2022) ($\alpha=256$, $\text{rank}=128$, $\text{dropout}=0.05$) using 8-bit quantization. We use the AdamW Loshchilov and Hutter (2019) optimizer for LLM and $W_{m}$ parameter with a batch size of $32$ and RMSprop optimizer for the LSTM and perform a grid search to determine the optimal learning rate. In KCGen-KT, we set different learning rates for different model components: $51e-$ 5 for Llama 3, $45e-$ 4 for the LSTM model, and $41e-$ 4 for the $W_{m}$ and $b_{m}$ parameters. KCGen-KT converges within $12$ training epochs, with each epoch taking $80$ minutes on an NVIDIA L40S 48GB GPU. We repeat all our experiments with 5 random train-validation-test data splits for cross-validation.
For the baseline model TIKTOC* which also uses a LLM as the backbone of the model. We use same setup for a fair comparison: we use same Llama model with PEFT to load it and fine-tune it with LoRA ($\alpha=256$, $\text{rank}=128$, $\text{dropout}=0.05$) using 8-bit quantization. The optimizers and learning rate used are the same as KCGen-KT except the learning rate for LSTM model is $41e-$ 4.
For metrics, we use the rou_auc_score and f1_score from scikit-learn library to compute AUC and F1 metrics, respectively. For codeBLEU we adopt the official implementation provided in the microsoft/CodeXGLUE. In addition, we use scipy library to perform the hierarchical agglomerative clustering. All software we use in the development of this work is open source. We are consistent with the terms and intended use of all software and with OpenAI API.
Model AUC F1 Acc CodeBLEU
KCGen-KT 0.812 0.723 0.724 0.569
- Student submission = 0 0.789 0.674 0.704 0.529
- Student Code $\rightarrow$ AST 0.794 0.691 0.715 0.546
- Student Code $\rightarrow$ Generated 0.810 0.706 0.731 0.557
Table 5: Ablation study of KCGen-KT
Appendix C Ablation Study
Table 5 shows the results of the ablation study among all components of KCGen-KT on the dataset. For ablation conditions requiring code submissions for KC generation, we report results using two code solutions since we note that LLM-generated code tends to follow similar problem-solving strategies; we adopt the lowest level of KC abstraction and manually verify the correctness of all generated solutions. We see that including correct submissions is crucial; removing it results in performance decrease on both KT tasks, which suggests that it is difficult to capture all necessary KCs from just the problem statement alone. Using LLM-generated submissions instead of actual student submissions also results in a minor performance decrease, which suggests that actual student code is highly diverse and captures a more complete set of skills required in each problem. We also see that switching submissions to the abstract syntax tree (AST) representation as input to the LLM decreases performance, where we see the generated KCs tend to be less detailed. This result can be explained by LLMs not being heavily pre-trained on AST representations of code. See Appendix E for prompt used for KC generation with AST representation.
Furthermore, we explore the impact of the number of student submissions provided to the LLM on the quality of generated KCs and KT performance. For this ablation, we use the lowest KC abstraction level. Table 6 shows the results, where we see that using fewer submissions, such as 1 and 2, results in worse performance, while the performance increases and saturates after more than 5 submissions. This result can be explained by a smaller number of student submissions failing to capture diverse solution strategies for some problems, thus resulting in an incomplete KC set. As the number of submissions increase, the set of initial KCs before clustering more or less stays the same, and the performance mostly depends on the abstraction level instead.
No. of Solutions AUC F1 Acc CodeBLEU
1 0.804 0.713 0.705 0.563
2 0.812 0.723 0.724 0.569
5 0.815 0.726 0.737 0.572
7 0.812 0.726 0.727 0.573
10 0.816 0.727 0.715 0.566
Table 6: Ablation study of KCGen-KT on different No. of solution provided during KC Generation. Best performance is in bold and second best is underlined.
Appendix D Human Evaluation Details
@@
We conduct a human evaluation to assess the quality of generated KCs and problem-KC mapping accuracy. Our evaluators were volunteers contacted through a research partner and were not compensated monetarily. They were made aware that their annotations would be used in scientific research in AI. Annotators perform two separate tasks, each designed to evaluate a distinct aspect of the KC quality. Below, we describe the annotation rubrics and summarize the evaluation outcomes.
D.1 KC-Set Annotation Per Problem for Problem-KC mapping
@@
Which KC-set is better (1 for Generated, 2 for Baseline)? A KC-set is considered better if, overall, it more effectively captures the key concepts that are practiced in the problem and necessary for solving it. Imagine you are an instructor selecting problems for a homework assignment. Which set of concepts would be more informative in deciding whether to include this problem based on the skills it assesses?
To assess inter-rater reliability, we calculate Kappa scores. The overall agreement across all five annotators is relatively low, with an average Kappa score of 0.121. The highest pairwise agreement observed is 0.51, indicating moderate agreement between two annotators. This disagreement suggests differing interpretations among annotators regarding which KCs are most relevant to a given problem. In particular, the task excludes KC understandability as a criterion, focusing instead on relevance and completeness. These differences in evaluative emphasis and background knowledge likely contributed to the observed disagreement.
D.2 Single-KC Annotation Per Problem for Generated KC Quality
In the second task, annotators evaluate the quality of individual KCs generated for each problem by assessing their correctness and instructional relevance. This task focuses on the Problem-KC mappings accuracy.
The annotation rubric is as follows:
Task: For each KC associated with a problem, provide a binary label for correctness, and if the KC is correct, further classify its level of essentiality.
Correct: if the KC accurately reflects a concept required to understand or solve the problem, and it holds educational value in the context.
Essentiality (only if the KC is Correct): Classify the KC based on its instructional importance using the following three categories:
  • Essential: The KC is critical to the problem. It represents a core concept that the problem is explicitly designed to assess. Without understanding this concept, a student would likely struggle to solve the problem.
  • Marginal: The KC is relevant but not central. It may influence the decision to assign the problem, but only as a secondary consideration. While the problem can help reinforce this concept, it is not the primary instructional goal.
  • Non-Essential: The KC appears in the problem but plays a minimal instructional role. It is not a focus of the problem and would not factor significantly into an instructor’s decision to use the problem for teaching purposes.
Figure 5: A demo of the interface for human evaluation
Results show that the average percentage of relevant KCs per problem is $92.0\%$ for the LLM-generated set and $91.6\%$ for the baseline. The average percentage of essential KCs per problem is $50.1\%$ for the generated set and $49.5\%$ for the baseline, the average percentage of marginal KCs is $31.9\%$ for generated KCs and $33.7\%$ for the baseline KCs and the average percentage of non-essential KCs is $31.9\%$ and $33.7\%$, respectively. These findings suggest that the LLM-based KC tagging is reasonably accurate and LLM-generated KCs are largely relevant and instructionally meaningful, which implies that LLM is capable of identifying topic-specific concepts with comparable instructional value, but still have room to improve with human-AI collaboration approach.
Annotators complete the evaluation using a custom web-based interface, as shown in figure 5,that presented each problem alongside its associated KC set. For each KC, annotators labeled correctness and essentiality and selected their preferred KC set. The interface was designed to ensure consistency and ease of use across annotators.
Appendix E Prompt
E.1 Prompt for KC Generation Pipeline
We show the prompt used for the KC generation for each problem in Table E.3, the prompt used for the AST ablation study in Table E.3, and the prompt used for cluster summarization in Table 10.
E.2 Prompt for KC Correctness Labeling
We show the prompt used for KC correctness labeling for the learning curve analysis in Table E.3.
E.3 Prompt for In-context Example Conversion

### v3

@@ 5.2.1. Case Study
Table 5.1.3 shows the estimated KC mastery levels and predicted code submission for a student on a problem from the CodeWorkout test set. The low student mastery level on KCs, “For loop iteration” and “Array indexing and assignment”, results in a run-time error in the predicted code by indexing the array outside of its bounds. In contrast, the higher predicted mastery level of other KCs results in the correct implementation of the if and else if statements, proper use of boolean expressions, and accurate application of the logical AND operator. This example shows that informative KC descriptions generated by the LLM can help KCGen-KT make more accurate student code predictions. In practice, the predicted mastery level may offer instructors interpretable insights into the student’s understanding of specific programming concepts.
5.2.2. Learning Curve Analysis
@@
A common method to assess the quality of KCs is examining how well they match cognitive theory; the expected pattern on the KCs should follow the power law of practice, which states that the number of errors should decrease as the amount of practice on certain KCs increases (Newell and Rosenbloom, 2013; Snoddy, 1926). Hence, we compare the error rate across different attempts at KCs with the estimated student KC mastery levels from KCGen-KT on CodeWorkout. For this experiment, we prompt GPT-4o to label whether each student submission contains an error on each KC. For all incorrect student submissions, we provide the problem statement, the associated KCs, and the student code, and prompt GPT-4o to (1) reason about the errors, (2) generate a corrected version of the code, and (3) assign a binary correctness label to each KC, indicating whether the student made an error on this KC in their submission (See Appendix B for the exact prompt). We acknowledge that although we find this process to be empirically accurate, a formal validation is necessary to verify the accuracy of the KC-level correctness labels.
@@
Figure 3. Representative learning curves for three generated KCs (Equality Comparison, String Length Determination, and For Loop Iteration), showing a generally decreasing error rate over attempts. Our KCs result in better model fit (0.21 vs. 0.18 in $R^{2}$) than human-written KCs under cognitive models (Pavlik et al., 2009).:
@@
Figure 3 shows three representative learning curves among all LLM-generated KCs. In all cases, both the ground truth and predicted error rate curves exhibit a general decreasing trend as the number of attempts increases, consistent with the power law of practice. The first predicted learning curve closely aligns with the ground truth, demonstrating KCGen-KT’s ability to accurately capture student learning progressions. The second predicted learning curve matches the ground truth error rates in trend, but has higher overall values. The third curve further exacerbates this discrepancy, for a KC that appears more frequently in the dataset. The ground truth error rates decrease overall but have significant fluctuation, making it difficult to fit the predicted learning curve. The likely reason is that students attempt problems in different orders in the CodeWorkout dataset, with some students skipping certain questions; this variation means that the same attempt may correspond to questions with different difficulties across students, making the average error rate noisy.
For a more quantitative evaluation, following prior work (Pavlik et al., 2009), we fit PFA models on each KC. LLM-generated KCs achieve a weighted $R^{2}$ of $0.21$ versus $0.18$ using human-written KCs on CodeWorkout. Therefore, the LLM-generated KCs fit the power law of practice slightly better compared to human-written KCs.
5.2.3. KC Ontologies
@@
We show a portion of the CodeWorkout KC ontology subtree obtained from KCGen-KT, for string-related concepts, in Figure 4. The root node shows the KC at the highest abstraction level; going down the tree from there, we see how the KCs identified increase in granularity at each level. To build the ontology tree, we start from the top-level KC and identify all KC labels from the next level that are semantically included in it, and do it iteratively for all KCs. This mapping defines the parent-child relationships between KCs across different abstraction levels.
This example demonstrates the controllable abstraction property of our KC generation pipeline, where adjusting the number of clusters directly controls the granularity of the generated KCs.
@@
Figure 4. A section of the generated KC ontology (related to Strings), at different abstraction levels.:
6. Human Evaluation
We perform a human evaluation to assess the quality of the generated KCs and the accuracy of problem-KC mappings. We recruit two annotators with experience in teaching college-level programming in Java. Since their expertise lies in Java-based instruction, we focus the evaluation on all $50$ CodeWorkout problems and provide them three separate tasks, each targeting a distinct aspect.
To evaluate the interpretability of KCs, we provide the generated KCs for each problem and ask annotators to indicate whether each KC is clearly understandable from an instructor’s perspective. On average, $94.6\%$ of baseline KCs and $98.6\%$ of LLM-generated KCs are labeled as interpretable. Inter-annotator agreement is adequate: for example, the two human annotators reach a Cohen’s Kappa score of 0.594 when labeling LLM-generated KCs, indicating moderate (on the borderline between moderate and good) agreement. This result suggests that the quality of generated KCs is comparable to or more informative than the human-created baseline KCs, highlighting their potential for instructional use in practice.
To evaluate problem-KC mappings, we evaluate the precision and recall aspects, measuring whether labeled KCs accurately and completely capture the skills required by each problem. Since there is no “ground-truth”, for precision, we ask annotators to label each KC as relevant or irrelevant to the problem. For recall, we ask them to perform a head-to-head comparison between two sets of KCs and rate, in binary format, whether the LLM-generated set has less coverage or equal-or-greater coverage of essential KCs relative to the baseline. We then compute the average proportion of relevant KCs per problem and the proportion of cases where the generated set achieves equal or greater coverage.
Results show that the average precision is $92.5\%$ for the baseline set and is $93.2\%$ for the LLM-generated set with an inter-annotator agreement of $89.3\%$. For recall, the LLM-generated set is judged to have equal or higher coverage in $96\%$ of the cases, with $92\%$ agreement. We note that surpassing the baseline human-authored KC set on these tasks is highly challenging, making these results particularly encouraging. Overall, these findings indicate that LLM-based KCs are both interpretable and comparable to human-created baselines. See Appendix A for the detailed annotation rubrics.
7. Conclusions and Future Work
@@
In this paper, we presented an automated, LLM-based pipeline for KC generation and tagging for open-ended programming problems. We also developed an LLM-based KT framework, KCGen-KT, which leverages the textual content of KC descriptions. Through extensive experiments on two real-world coding datasets, we showed that KCGen-KT outperforms human-written KC labels on KC generation and existing state-of-the-art KT methods on predicting future student performance. We also show that LLM-generated KCs lead to better learning curves than human-written ones. A human evaluation shows that the generated KCs and problem-KC mappings are reasonably accurate for programming instructors.
@@
There are many avenues for future work. First, a human-in-the-loop KC generation method may address the limitation that KCGen-KT sometimes misses KCs shared across problems due to generating KCs independently for each problem. Second, we can explore integrating the learning curve fit directly into the objective to explicitly optimize for the power law of practice. Third, we plan to investigate applying our approach to other student modeling domains such as dialogues (Scarlatos et al., 2025), math (Ozyurt et al., 2024), and science (Moore et al., 2024).
Appendix A Human Evaluation Details
@@
We conduct a human evaluation to assess (1) the interpretability of generated KCs as a measure of their instructional quality and (2) the precision and recall of the problem–KC mappings. Our evaluators were volunteers contacted through a research partner and were not compensated monetarily. They were made aware that their annotations would be used in scientific research in AI. The detailed annotation rubrics are provided below.
A.1. KC Interpretability
@@
For each KC, indicate whether it is clearly understandable from an instructor’s perspective to decide the KC’s interpretability using Yes or No.
A.2. Problem KC Mapping
We consider the correctness of problem-KC mapping from both precision and recall perspective. Precision measures whether KCs assigned to a problem are relevant. Recall metric measures whether the KC set sufficiently covers the essential concepts required for solving the problem. Because no absolute ground-truth KC set exists for each problem, recall is measured using relative coverage between the baseline and LLM-generated KC sets. The instruction provided to evaluate the problem-KC mapping is as follow:
Precision: For each KC associated with a problem, indicate whether it is relevant to that problem in binary format (relevant/irrelevant). Recall: For each problem, compare the generated KC set and baseline KC set and assign a binary label indicating which set better captures the essential KCs. Label ”Less coverage” if Set 1 covers fewer relevant KCs than Set 2 and label ”Equal-or-greater coverage” if Set 1 covers more relevant KCs than Set 2. If both sets contain unique KCs that the other is missing which means neither list clearly dominates, select equal-or-greater coverage.
Appendix B Prompt
B.1. Prompt for KC Generation Pipeline
We show the prompt used for the KC generation in Table LABEL:tab:kc_generation_prompt and the prompt used for cluster summarization in Table LABEL:tab:kc_summarize_prompt.
B.2. Prompt for KC Correctness Labeling
We show the prompt used for KC correctness labeling for the learning curve analysis in Table LABEL:tab:kc_correctness_label.
B.3. Prompt for In-context Example Conversion